the following hypothesis are given

H0:p=0

h1:p#0

A random sample of 12 paired observations indicated a correlation of 0.32. Can we

conclude that the correlation in the population is equal to zero? Use 0.05 level of

significance. (3 marks)

the following sample observations were randomly selected

x 4 5 3 6 10

y 4 6 5 7 7


1
Expert's answer
2021-11-15T17:59:23-0500

a.

From the data above, the correlation coefficient is "r=0.32" and "n=12"

The hypothesis tested are,

"H_0:\rho=0\space against \space H_1:\rho\not=0"

and apply the student's t distribution to make a decision.

The test statistic is given as,

"t^*=r\sqrt{n-2}/\sqrt{1-r^2}=0.32\sqrt{10}/\sqrt{1-(0.32)^2}= 1.068092"

"t^*" is compared with the table value at "\alpha =0.05" with "n-2=12-2=10" degrees of freedom.

Now, "t_{\alpha/2,10}=t_{0.05/2,10}=t_{0.025,10}=2.228" and the null hypothesis is rejected if "|t^*|\gt t_{0.025,10}"

Since "|t^*|=1.068092\lt t_{0.025,10}=2.228", we fail to reject the null hypothesis and conclude that there is sufficient evidence to show that the correlation coefficient of "r=0.32"  in the population is equal to zero at 5% level of significance.


b.

To find the regression equation, we enter the following commands in "R".

x=c(4,5,3,6,10)

y=c(4,6,5,7,7)

lm(y~x)

The output for these commands is

Call:

lm(formula = y ~ x)


Coefficients:

(Intercept)      x  

   3.767    0.363 

This shows that the "y-intercept=3.767" and the "slope=0.363". The regression line is of the form "y=a*x+b" where "a" is the slope and "b" is the y-intercept. Therefore, the regression line equation is, "y=0.363x+3.767"

When x=7, the value of y is obtained by substituting for x in the regression line equation line as follows,

y=(0.363*7)+3.767= 6.308

Therefore, the value of y when x=7 is 6.308.